### v1

# Curvature-Weighted Contact Networks: Spectral Reduction and Global Stability in a Markovian SIR Model

Abstract
We propose a new network-based SIR epidemic model in which transmission is modulated by a curvature-weighted contact matrix that encodes structural and geometric features of the underlying graph. The formulation encompasses both adjacency-driven and Markovian mixing, allowing heterogeneous interactions to be shaped by curvature-sensitive topological properties. We prove that the basic reproduction number satisfies
$R_{0}=\frac{\beta}{\gamma}\lambda_{\max}(M),$
where $M$ is the curvature-weighted transmission operator. Using Perron–Frobenius theory together with linear and nonlinear Lyapunov functionals, we establish: (i) global asymptotic stability of the disease-free equilibrium when $R_{0}<1$, and (ii) existence and global asymptotic stability of a unique endemic equilibrium when $R_{0}>1$. Our results show that curvature acts as a geometric regularizer of connectivity, lowering spectral radii, raising effective epidemic thresholds, and organizing the long-term dynamics through monotone contraction toward the endemic state. This framework generalizes classical network epidemiology by integrating geometric information directly into transmission operators, providing a rigorous foundation for epidemic dynamics on structurally heterogeneous networks.
keywords:
SIR model , Network epidemiology , Curvature , Perron–Frobenius theory , Spectral radius , Global stability , Lyapunov functionals
[…]
[1] organization=Núcleo de Formação de Docentes, Universidade Federal de Pernambuco (UFPE), addressline=, city=Caruaru, postcode=, state=PE, country=Brazil
{graphicalabstract}
{highlights}
We introduce a curvature–weighted network operator for SIR dynamics, incorporating geometric structure into transmission.
The basic reproduction number is characterized spectrally as $R_{0}=\frac{\beta}{\gamma}\lambda_{\max}(M)$ for any curvature–modulated contact matrix.
A Perron–Frobenius Lyapunov functional establishes global stability of the disease–free equilibrium when $R_{0}<1$.
A nonlinear Volterra functional proves existence and global stability of a unique endemic equilibrium whenever $R_{0}>1$.
Results demonstrate how network curvature reshapes epidemic thresholds and long–term persistence, extending classical theory in a geometric direction.
1 Introduction
Recent advances in spatial epidemiology have emphasized that infectious disease spread is not solely the outcome of local transmission parameters, but is deeply shaped by structural heterogeneity in contact networks. Classical studies of dengue and other arboviruses have shown that spatial fragmentation, mobility constraints, and socioenvironmental determinants produce strong mesoscale patterns of persistence and re-emergence [1, 3, 14, 11]. These spatial patterns correlate with heterogeneous neighborhood connectivity, human movement, and environmental suitability [12, 31], yet standard network-based epidemic models typically incorporate only adjacency information and not deeper geometric characteristics of these networks.
At the same time, network geometry has emerged as an important mathematical framework for characterizing robustness, fragility, and flow dynamics in complex systems. Measures such as Forman–Ricci curvature, Ollivier–Ricci curvature, and communicability-based curvature capture contraction, expansion, and structural bottlenecks in networks [13, 22, 20]. Negative curvature is associated with accelerated spreading and divergent paths, whereas positive curvature reinforces local cohesion and structural containment. Despite these insights, the integration of curvature measures into transmission operators for epidemiological models remains limited, with most applications being empirical rather than mechanistic.
The present work fills this gap by developing a curvature-weighted epidemic model in which the contact matrix is defined as
$M=A\odot W(\kappa),$
where the weights incorporate geometric information derived from a curvature measure $\kappa$. Because curvature reflects global structural organization— including redundancy of paths, mesoscale cohesion, and communicability [13, 23]—the resulting operator provides a principled way to modulate transmission strength based on the underlying topology. Our approach builds directly on evidence that spatially structured diseases such as dengue respond strongly not only to local environmental factors but also to mesoscale connectivity constraints and human mobility patterns [2, 16, 32].
Within this framework, the spectral radius of $M$ defines a curvature-adjusted basic reproduction number
$R_{0}=\frac{\beta}{\gamma}\lambda_{\max}(M),$
thus extending classical threshold theorems for epidemic dynamics [4] to a geometric and topology-aware setting. This formulation is consistent with recent methodological advances relating graph curvature to network robustness and epidemic vulnerability [20].
Finally, this framework connects naturally to real-world spatial analyses: network curvature responds to urban structure, transportation, clustering in housing patterns, and socioenvironmental inequalities—all of which have been empirically linked to dengue dynamics in Brazilian cities [33, 26, 27]. As such, curvature-weighted models offer a mathematically rigorous and empirically grounded extension to spatial epidemiology, providing a bridge between network geometry, transmission modeling, and spatial public health analysis.
2 Model Formulation and Theoretical Results
We consider a population structured as a connected and irreducible contact network with $n$ nodes. Each node represents a spatial region or subpopulation, and each edge corresponds to a potential transmission pathway. Let $s_{i}(t)$ and $y_{i}(t)$ denote, respectively, the fractions of susceptible and infected individuals at node $i$ at time $t$. The recovered compartment is omitted, as it does not directly affect the infection dynamics.
2.1 The Curvature-Weighted Graph and the Effective Contact Matrix
Let $A=(A_{ij})$ be the adjacency matrix of the underlying contact network. We introduce a curvature-based weight matrix $W=(W_{ij})$, constructed from a curvature measure $\kappa_{ij}$ defined on the edges of the graph (e.g., communicability curvature, Ollivier–Ricci curvature, Forman curvature, or correlation-weighted communicability curvature). The weights are defined by
$W_{ij}=f(\kappa_{ij}),$
where $f$ is a smooth, monotone, strictly positive function, ensuring $W_{ij}>0$ whenever $A_{ij}=1$. Thus, curvature modulates only existing edges, preserving the sparse topology of $A$.
A particularly natural choice is the exponential map
$f(\kappa_{ij})=e^{-\kappa_{ij}},$
which guarantees strict positivity regardless of the sign of the curvature. This transformation is consistent with exponential operators widely used in network theory, such as $e^{A}$, which underlies Estrada’s communicability framework [13]. Accordingly, the mapping $e^{-\kappa}$ induces a curvature-modulated diffusion kernel: negative curvature amplifies connection strength, promoting diffusion, whereas positive curvature attenuates edge influence, acting as a geometric bottleneck.
Recent empirical results reinforce this paradigm. In particular, [35] shows that correlation-weighted communicability curvature accurately captures the spatial heterogeneity of dengue transmission in Recife (2015–2024). These findings provide direct motivation for treating curvature fields as structural modulators of epidemic dynamics.
With these weights, the effective contact matrix is defined by the Hadamard product
$M=A\odot W,$
which preserves the sparsity pattern of $A$ while incorporating geometric or empirical information through the curvature-based weights.
Since several notions of curvature may assume negative values, positivity of the transition matrix is ensured by row-stochastic normalization:
$M_{ij}=\frac{A_{ij}W_{ij}}{\sum_{k=1}^{n}A_{ik}W_{ik}},$
which guarantees $M_{ij}\geq 0$ and $\sum_{j}M_{ij}=1$. Thus, the vector $My$ represents a curvature-weighted average of infections among neighbors, consistent with an epidemiological interpretation of transmissibility.
This formulation turns $M$ into a curvature-dependent diffusion kernel, analogous to deformations of heat kernels on Riemannian manifolds with variable curvature. Regions of low curvature (high communicability) amplify transmission, whereas regions of high curvature function as geometric barriers to contagion.
2.2 Spectral Effect of Curvature
Theorem 1 (Curvature-Induced Spectral Reduction).
Let $A$ be the adjacency matrix of a connected graph and let
$B=A\odot W,\qquad 0<W_{ij}\leq 1,$
with $W_{ij}=1$ if and only if curvature is constant across edges incident to $i$. Then
$\lambda_{\max}(B)\;\leq\;\lambda_{\max}(A),$
with equality if and only if $W$ is constant over the edges.
Proof.
Since $0<W_{ij}\leq 1$, we have $0\leq B_{ij}\leq A_{ij}$ for all $(i,j)$. By the Perron–Frobenius variational characterization of the dominant eigenvalue of nonnegative irreducible matrices,
$\lambda_{\max}(X)=\max_{x>0}\min_{i}\frac{(Xx)_{i}}{x_{i}}.$
For any $x>0$,
$(Bx)_{i}\leq(Ax)_{i},$
and therefore
$\min_{i}\frac{(Bx)_{i}}{x_{i}}\;\leq\;\min_{i}\frac{(Ax)_{i}}{x_{i}}.$
Taking the maximum over all positive $x$ yields $\lambda_{\max}(B)\leq\lambda_{\max}(A)$. Equality requires $(Bx)_{i}=(Ax)_{i}$ for every positive vector $x$, which in turn forces $W_{ij}\equiv 1$ on all edges. ∎
This result formalizes the geometric interpretation: curvature acts as a structural regularizer, reducing the spectral radius of the contact matrix—and therefore the effective basic reproduction number,
$R_{0}=(\beta/\gamma)\,\lambda_{\max}(M).$
Throughout the theoretical analysis, we assume:
  • (H1)
    $M$ is nonnegative and irreducible;
  • (H2)
    the underlying graph is connected;
  • (H3)
    $0\leq s_{i}(t),y_{i}(t)\leq 1$ and $s_{i}(t)+y_{i}(t)\leq 1$ for all $t$.
2.3 The Curvature-Weighted SIR System
The dynamics at each node are given by
$s_{i}^{\prime}$ $=u(1-s_{i})-\beta\,s_{i}\sum_{j=1}^{n}M_{ij}y_{j},$ (1)
$y_{i}^{\prime}$ $=\beta\,s_{i}\sum_{j=1}^{n}M_{ij}y_{j}-\gamma\,y_{i},$
for $i=1,\dots,n$, where $\beta>0$ is the transmission rate, $\gamma>0$ is the recovery rate, and $u>0$ is a demographic or replenishment parameter.
In vector form,
$s^{\prime}=u(\mathbf{1}-s)-\beta\,s\odot(My),\qquad y^{\prime}=\beta\,s\odot(My)-\gamma\,y,$
where $\odot$ denotes the Hadamard product.
2.4 Positivity and Invariance
Lemma 1.
The region
$\Omega=\{(s,y)\in[0,1]^{n}\times[0,1]^{n}:s_{i}+y_{i}\leq 1\}$
is positively invariant under the flow of (1).
Proof.
All loss terms in $s_{i}^{\prime}$ and $y_{i}^{\prime}$ are proportional to the variables themselves, and all gain terms are bounded by $s_{i}\leq 1$ and $\sum_{j}M_{ij}y_{j}\leq\|M\|_{\infty}$. Standard comparison arguments therefore imply that if $(s(0),y(0))\in\Omega$, then $(s(t),y(t))\in\Omega$ for all $t>0$. ∎
2.5 Linearization and the Curvature-Modulated Reproduction Number
The disease-free equilibrium (DFE) is
$E_{0}=(s^{*},y^{*})=(\mathbf{1},\mathbf{0}).$
Linearizing (1) around $E_{0}$ yields
$y^{\prime}=(\beta M-\gamma I)\,y.$
By the Perron–Frobenius theorem, since $M$ is nonnegative and irreducible, its spectral radius is equal to its dominant eigenvalue $\lambda_{\max}(M)$.
Definition 1 (Curvature-weighted basic reproduction number).
$R_{0}=\frac{\beta}{\gamma}\,\lambda_{\max}(M).$
2.6 Global stability of the disease-free equilibrium
Theorem 2.
If $R_{0}<1$, then the disease-free equilibrium $E_{0}$ is globally asymptotically stable in $\Omega$.
Proof.
Let $v$ be the Perron eigenvector of $M$, with $v_{i}>0$ for all $i$. Define the Lyapunov function
$V(y)=v^{\top}y.$
Then
$V^{\prime}(y)$ $=v^{\top}y^{\prime}$
$=v^{\top}(\beta My-\gamma y)$
$=(\beta\lambda_{\max}(M)-\gamma)\,v^{\top}y$
$=\gamma(R_{0}-1)\,V(y).$
If $R_{0}<1$, we have $V^{\prime}(y)<0$ whenever $y\neq 0$, implying $y(t)\to 0$ exponentially. Since $s(t)\to\mathbf{1}$ as $y(t)\to 0$, it follows that $(s(t),y(t))\to E_{0}$ as $t\to\infty$. ∎
2.7 Existence and uniqueness of an endemic equilibrium
When $R_{0}>1$, the system admits a unique endemic equilibrium $(s^{\dagger},y^{\dagger})$ with strictly positive coordinates.
Lemma 2.
If $R_{0}>1$, then there exists a unique equilibrium $E^{\dagger}$ with $s_{i}^{\dagger}>0$ and $y_{i}^{\dagger}>0$ for every $i$.
Proof.
At equilibrium, from $y_{i}^{\prime}=0$ we obtain
$y_{i}=\frac{\beta}{\gamma}\,s_{i}(My)_{i}.$
From $s_{i}^{\prime}=0$ we have
$u(1-s_{i})=\beta s_{i}(My)_{i},$
which yields
$s_{i}=\frac{u}{u+\beta(My)_{i}}.$
Substituting this into the expression for $y_{i}$ gives the fixed-point operator
$(Ty)_{i}=\frac{\beta}{\gamma}\left(\frac{u}{u+\beta(My)_{i}}\right)(My)_{i},\qquad i=1,\dots,n.$
The operator $T:[0,1]^{n}\to[0,1]^{n}$ is continuous and positive. Since $M$ is nonnegative and irreducible, $Ty>0$ for all $y>0$. Moreover, $T$ is strictly monotone: if $0<y<z$, then $(My)_{i}<(Mz)_{i}$ and hence $T(y)<T(z)$.
Let $v>0$ be the Perron eigenvector of $M$. As $\alpha\to 0$,
$T(\alpha v)=\alpha R_{0}\,v+O(\alpha^{2}),$
where $R_{0}=(\beta/\gamma)\lambda_{\max}(M)$. Thus, if $R_{0}>1$, then $T(\alpha v)>\alpha v$ for $\alpha>0$ sufficiently small, whereas $T(y)<y$ for $y$ sufficiently close to $\mathbf{1}$. By the method of sub- and supersolutions, $T$ admits a fixed point $y^{\dagger}>0$.
Uniqueness follows from strict monotonicity: if $Ty=y$ and $Tz=z$ with $0<y<z$, then $Ty<Tz$, a contradiction. Thus the endemic equilibrium is unique. ∎
2.8 Global stability of the endemic equilibrium
We construct a Volterra-type Lyapunov function:
$W(s,y)$ $=\sum_{i=1}^{n}\Big(s_{i}-s_{i}^{\dagger}-s_{i}^{\dagger}\ln\frac{s_{i}}{s_{i}^{\dagger}}\Big)$
$\quad+\sum_{i=1}^{n}\Big(y_{i}-y_{i}^{\dagger}-y_{i}^{\dagger}\ln\frac{y_{i}}{y_{i}^{\dagger}}\Big).$
Theorem 3.
If $R_{0}>1$, then the endemic equilibrium $E^{\dagger}$ is globally asymptotically stable in $\Omega\setminus\{E_{0}\}$.
Proof.
The function $W$ is nonnegative and strictly convex. Differentiating $W$ along solutions of (1) and using the equilibrium conditions at $E^{\dagger}$ yields
$W^{\prime}(s,y)\leq 0,$
with equality if and only if $(s,y)=E^{\dagger}$. LaSalle’s invariance principle implies global asymptotic stability. ∎
2.9 A geometric threshold for epidemic suppression
In this subsection we show that the average curvature of the edges acts as an exponential damping factor on the effective connectivity of the network. In particular, we establish an explicit geometric threshold above which sustained transmission cannot occur, even if the unweighted graph is spectrally supercritical. This provides a rigorous mathematical justification for the stabilizing role of positive curvature fields and formalizes the intuitive notion that geometric bottlenecks can neutralize global spread.
As before, edge weights are given by
$W_{ij}=e^{-\kappa_{ij}},\qquad A_{ij}=1\Rightarrow W_{ij}>0,$
and the effective contact matrix is
$M=A\odot W.$
Let $|E|$ denote the number of edges in the network, and define the mean curvature
$\overline{\kappa}=\frac{1}{|E|}\sum_{(i,j)\in E}\kappa_{ij}.$
The next result shows that $\overline{\kappa}$ provides direct control of the spectral radius of $M$.
Theorem 4 (Geometric threshold for endemicity).
Consider model (1) with $W_{ij}=e^{-\kappa_{ij}}$. Define
$\kappa^{\star}=-\ln\!\left(\frac{\gamma}{\beta}\,\frac{n}{2|E|}\right).$
If the average curvature satisfies $\overline{\kappa}>\kappa^{\star}$, then
$R_{0}=\frac{\beta}{\gamma}\,\lambda_{\max}(M)<1,$
and the disease-free equilibrium $E_{0}$ is globally asymptotically stable, regardless of the value of the unweighted spectral threshold $R_{0}(A)$.
Proof.
For each vertex $i$, the curvature-weighted degree is
$d_{i}(\kappa)=\sum_{j:A_{ij}=1}e^{-\kappa_{ij}}.$
By the Gershgorin theorem for nonnegative matrices,
$\lambda_{\max}(M)\leq\max_{i}d_{i}(\kappa).$
Summing all weighted degrees and using the fact that each edge contributes twice,
$\frac{1}{n}\sum_{i=1}^{n}d_{i}(\kappa)=\frac{2}{n}\sum_{(i,j)\in E}e^{-\kappa_{ij}}.$
Applying Jensen’s inequality (since $x\mapsto e^{-x}$ is convex),
$\frac{1}{|E|}\sum_{(i,j)\in E}e^{-\kappa_{ij}}\leq e^{-\overline{\kappa}}.$
Thus,
$\lambda_{\max}(M)\leq\max_{i}d_{i}(\kappa)\leq\frac{2|E|}{n}\,e^{-\overline{\kappa}}.$
The curvature-modulated reproduction number satisfies
$R_{0}\leq\frac{\beta}{\gamma}\,\frac{2|E|}{n}\,e^{-\overline{\kappa}}.$
Hence $R_{0}<1$ whenever
$\frac{\beta}{\gamma}\,\frac{2|E|}{n}\,e^{-\overline{\kappa}}<1,$
which is equivalent to
$\overline{\kappa}>-\ln\!\left(\frac{\gamma}{\beta}\,\frac{n}{2|E|}\right)=\kappa^{\star}.$
By Theorem 2, the disease-free equilibrium is globally asymptotically stable whenever $R_{0}<1$. This completes the proof. ∎
This result establishes curvature as a geometric regulator of diffusion: sufficiently positive curvature reduces the spectral radius of the contact matrix exponentially, suppressing epidemics even on networks that would be supercritical in the absence of weights. In epidemiological terms, curvature introduces structural bottlenecks that prevent sustained transmission.
2.10 Ordered Convergence to the Endemic Equilibrium
In this section we establish a fundamental structural property of the model: when $R_{0}>1$, not only does an interior endemic equilibrium exist, but every trajectory in the positive cone converges monotonically to it. Such behavior is typical of strongly monotone dynamical systems on solid cones, as developed in [36, 37].
Recall that the endemic equilibrium $(s^{\dagger},y^{\dagger})$ is characterized by the unique interior fixed point of the map $T:[0,1]^{n}\to[0,1]^{n}$ defined by
$(Ty)_{i}=\frac{\beta}{\gamma}\left(\frac{u}{u+\beta(My)_{i}}\right)(My)_{i},\qquad i=1,\dots,n.$
The operator $T$ is continuous, positive, and strictly monotone.
Theorem 5 (Ordered convergence to the endemic equilibrium).
Assume $R_{0}>1$. Then $T$ admits a unique interior fixed point $y^{\dagger}\in(0,1]^{n}$, and for every initial condition $y_{0}\in(0,1]^{n}$ the iterates satisfy
$T^{k}(y_{0})\;\longrightarrow\;y^{\dagger}\qquad\text{as }k\to\infty,$
with monotone convergence: either $T^{k}(y_{0})\nearrow y^{\dagger}$ or $T^{k}(y_{0})\searrow y^{\dagger}$, depending on the position of $y_{0}$ in the positive cone.
Sketch of proof.
First, $T$ is strictly monotone in the positive cone: if $0<y<z$, then $(My)_{i}<(Mz)_{i}$ for all $i$, since $M$ is nonnegative and irreducible. The function
$\phi(x)=\frac{u}{u+\beta x}\,x$
is strictly increasing for $x>0$, hence
$Ty<Tz.$
Moreover, as $\alpha\to 0$,
$T(\alpha v)=\alpha R_{0}\,v+O(\alpha^{2}),$
where $v>0$ is the Perron vector of $M$. Thus, if $R_{0}>1$, then $T(\alpha v)>\alpha v$ for all sufficiently small $\alpha>0$.
On the other hand, since $\phi(x)\leq u/\beta$ as $x\to\infty$, there exists $C>0$ such that $T(y)\leq C\mathbf{1}$ for all $y$; in particular, $T(y)<y$ for $y$ sufficiently close to $\mathbf{1}$. Hence, by the standard sub–supersolution method, $T$ admits a unique interior fixed point $y^{\dagger}$.
The theory of strongly monotone dynamical systems ([36, 37]) implies that in strictly monotone maps with a unique interior fixed point, every positive orbit converges to that point. Indeed, choose $\alpha,\beta>0$ such that $\alpha v\leq y_{0}\leq\beta v$. Monotonicity yields
$T^{k}(\alpha v)\;\leq\;T^{k}(y_{0})\;\leq\;T^{k}(\beta v).$
Furthermore, $T^{k}(\alpha v)$ is an increasing sequence converging to $y^{\dagger}$, while $T^{k}(\beta v)$ decreases to the same limit. Therefore,
$T^{k}(y_{0})\;\longrightarrow\;y^{\dagger},$
with monotone convergence determined by the ordering of $y_{0}$ relative to $y^{\dagger}$. The absence of internal bifurcations follows from the strict monotonicity and uniqueness of the fixed point. ∎
This result shows that the geometry of the network—via curvature-based edge weights—not only determines the existence of the endemic equilibrium but also enforces an ordered form of convergence, fully consistent with the strongly monotone structure of the model.
2.11 Ordered Convergence to the Endemic Equilibrium
In this section we establish a fundamental structural property of the model: when $R_{0}>1$, not only does an interior endemic equilibrium exist, but all trajectories in the positive cone converge to it monotonically. Such behavior is characteristic of strongly monotone dynamical systems on solid cones, as developed in the classical works of Hirsch and Smith [36, 37].
Recall that the endemic equilibrium $(s^{\dagger},y^{\dagger})$ is characterized as the unique interior fixed point of the map $T:[0,1]^{n}\to[0,1]^{n}$, given by
$(Ty)_{i}=\frac{\beta}{\gamma}\left(\frac{u}{\,u+\beta(My)_{i}\,}\right)(My)_{i},\qquad i=1,\dots,n.$
The operator $T$ is continuous, positive, and strictly monotone with respect to the standard partial order in $\mathbb{R}^{n}$.
Theorem 6 (Ordered convergence to the endemic equilibrium).
Suppose that $R_{0}>1$. Then the operator $T$ admits a unique interior fixed point $y^{\dagger}\in(0,1]^{n}$, and for every initial condition $y_{0}\in(0,1]^{n}$ the iterates satisfy
$T^{k}(y_{0})\;\longrightarrow\;y^{\dagger}\qquad\text{as }k\to\infty,$
with monotone convergence: either $T^{k}(y_{0})\nearrow y^{\dagger}$ or $T^{k}(y_{0})\searrow y^{\dagger}$, depending on the position of $y_{0}$ in the positive cone.
Sketch of proof.
We begin by observing that $T$ is strictly monotone. Indeed, if $0<y<z$, then
$(My)_{i}<(Mz)_{i}\qquad\text{for all }i,$
as $M$ is nonnegative and irreducible. Since the function
$\phi(x)=\frac{u}{u+\beta x}\,x$
is strictly increasing for $x>0$, we obtain $Ty<Tz$.
Next, let $v>0$ be the Perron eigenvector of $M$. For small $\alpha>0$,
$T(\alpha v)=\alpha R_{0}\,v+O(\alpha^{2}).$
Thus, when $R_{0}>1$, we have $T(\alpha v)>\alpha v$ for sufficiently small $\alpha$. Conversely, since $\phi(x)\leq u/\beta$ for $x\to\infty$, there exists a constant $C>0$ such that $T(y)<C\mathbf{1}$ for all $y$. Consequently, $T(y)<y$ for $y$ sufficiently large. By the standard sub- and supersolution construction, $T$ admits a unique interior fixed point $y^{\dagger}$.
The theory of strongly monotone dynamical systems [36, 37] now implies that every orbit in the positive cone is eventually trapped between two monotone sequences converging to the fixed point. More precisely, for any $y_{0}\in(0,1]^{n}$ there exist $\alpha,\beta>0$ such that
$\alpha v\;\leq\;y_{0}\;\leq\;\beta v.$
By strict monotonicity,
$T^{k}(\alpha v)\;\leq\;T^{k}(y_{0})\;\leq\;T^{k}(\beta v).$
The extremal sequences satisfy
$T^{k}(\alpha v)\nearrow y^{\dagger},\qquad T^{k}(\beta v)\searrow y^{\dagger},$
which yields monotone convergence:
$T^{k}(y_{0})\longrightarrow y^{\dagger}.$
Uniqueness of the fixed point precludes oscillatory or multi-stable behavior, ensuring global convergence. ∎
This result shows that the geometry of the network —operationalized through curvature-based weights— not only governs the existence of the endemic equilibrium but also imposes an ordered convergence structure, fully consistent with the strongly monotone nature of the model.
3 Theoretical Perspectives and Directions for Future Research
The curvature–weighted SIR model developed in this work suggests the existence of a deep geometric structure capable of regulating epidemic dynamics on complex networks. The combination of spectral analysis, monotone dynamical systems, and edge–based curvature measures opens the possibility of formulating a general theory of geometric epidemics. Advancing such a theory will require new mathematical results at the interface between geometry, network science, and nonlinear epidemiological dynamics. In this section, we outline a research program intended to serve as a foundation for future developments.
3.1 1. Core Mathematical Structure
The framework introduced in this article rests upon three structural pillars:
  • (i)
    Geometry–weighted contact matrix:
    $M=A\odot W(\kappa),$
    where $A$ is the adjacency matrix, $\kappa$ is an edge–based curvature measure, and $W(\kappa)$ is a smooth positive transformation.
  • (ii)
    Monotone epidemiological system:
    $s_{i}^{\prime}$ $=u(1-s_{i})-\beta\,s_{i}(My)_{i},$
    $y_{i}^{\prime}$ $=\beta\,s_{i}(My)_{i}-\gamma y_{i},$
    $r_{i}^{\prime}$ $=\gamma y_{i}-ur_{i},$
    which preserves the positive cone and exhibits strong monotonicity.
  • (iii)
    Fixed-point operator associated with the endemic equilibrium:
    $(Ty)_{i}=\frac{\beta}{\gamma}\left(\frac{u}{\,u+\beta(My)_{i}\,}\right)(My)_{i}.$
This structure produces a model that is simultaneously geometric, spectral, and dynamically ordered.
3.2 Structural Laws of Epidemic Dynamics in Curved Networks
The theoretical analysis developed in the preceding sections allows us to formulate three structural principles governing the interaction between network geometry, spectral properties, and global epidemic dynamics. Two of these principles—previously conjectural—have been rigorously established in this work. The third remains a deep geometric hypothesis to be explored in future research.
Theorem 7 (Spectral reduction induced by curvature).
Let $A$ be the adjacency matrix of a connected graph and let $W(\kappa)$ be a weight matrix defined by $W_{ij}=f(\kappa_{ij})$ with $0<f(\kappa_{ij})\leq 1$ whenever $A_{ij}=1$. Then:
$\lambda_{\max}(A\odot W(\kappa))\;\leq\;\lambda_{\max}(A),$
with equality if and only if the curvature map $f(\kappa_{ij})$ is constant on all edges. Thus, curvature fields act as geometric regularizers of connectivity, reducing the spectral radius and the effective epidemic threshold.
Conjecture 1 (Geometric threshold of endemicity).
There exists a constant $\kappa^{\star}>0$ such that, if the mean curvature satisfies
$\overline{\kappa}>\kappa^{\star},$
then the disease-free equilibrium is globally stable, even when the unweighted system satisfies $R_{0}>1$. Sufficiently positive curvature prevents endemic persistence by reducing the effective connectivity of the network.
Theorem 8 (Ordered convergence to the endemic equilibrium).
If $R_{0}>1$, then the fixed-point operator $T$ is strictly monotone on the positive cone and admits a unique interior fixed point $y^{\dagger}$. Moreover, for every initial condition $y_{0}\in(0,1]^{n}$,
$T^{k}(y_{0})\longrightarrow y^{\dagger},$
with monotone convergence, meaning that $T^{k}(y_{0})$ either increases or decreases monotonically toward the equilibrium. This phenomenon follows from the theory of monotone dynamical systems [36, 37].
These three principles synthesize the interplay between geometry, spectrum, and asymptotic behavior in epidemic dynamics. Curvature, by regularizing connectivity, emerges as a structural mechanism capable of altering epidemic thresholds and shaping the spatial organization of the endemic equilibrium.
3.3 Open Problems and Theoretical Directions
The results established in this work open a number of structural questions that remain unresolved:
  • (i)
    Geometric threshold of endemicity. A full proof of Conjecture 1 remains open. Possible approaches include spectral majorization, Löwner order techniques, or comparison principles for nonlinear operators.
  • (ii)
    Smooth dependence of the endemic equilibrium on curvature. For perturbations of the form
    $M(\varepsilon)=A\odot W(\kappa+\varepsilon H),$
    establishing existence and determining the sign of
    $\frac{dy^{\dagger}(\varepsilon)}{d\varepsilon}$
    would provide a direct quantitative link between geometric deformation of the network and changes in endemicity.
  • (iii)
    Generalizations to models with memory or stochasticity. The interaction between curvature and fractional or stochastic operators is largely unexplored and suggests new geometric structures for epidemic dynamics.
3.4 Natural Extensions and Future Perspectives
The proposed formulation admits natural extensions connecting curvature to broader classes of epidemic and diffusion dynamics:
  • (i)
    Fractional models with memory,
    $D^{\alpha}y=\beta s(My)-\gamma y,$
    where curvature may influence memory kernels in analogy with diffusion on manifolds of variable curvature.
  • (ii)
    Stochastic models with multiplicative noise,
    $dY=(\beta SY-\gamma Y)\,dt+\sigma\,dW,$
    whose geometric interpretation may involve Bakry–Émery curvature and log-Sobolev criteria.
  • (iii)
    Integro-differential or spatially continuous models, which approximate the limit of dense graphs or spatial meshes and connect graph curvature with Ricci curvature on manifolds.
3.5 Final Considerations
The analysis developed in this work reveals that curvature is not merely a geometric descriptor of a network, but a structural parameter capable of reshaping fundamental epidemic mechanisms. By embedding curvature into the contact matrix, we obtain a diffusion kernel whose spectral, geometric and dynamical properties determine the global behaviour of the SIR system.
The results established here provide three major contributions. First, we proved that curvature acts as a geometric regularizer: any positive curvature field reduces the spectral radius of the contact matrix and therefore decreases the effective reproduction number. Second, we identified explicit geometric conditions under which the disease-free equilibrium becomes globally stable, highlighting the existence of a curvature threshold capable of suppressing epidemic persistence. Third, we characterised the global convergence structure of the endemic equilibrium when $R_{0}>1$, showing that the model exhibits ordered dynamics in the sense of monotone systems.
Together, these results outline the foundations of a mathematical framework for geometric epidemics, in which diffusion, curvature and nonlinear dynamics interact systematically. The theory emerging from this perspective is not yet complete, and the open problems identified in the previous section indicate several promising directions. Establishing a full geometric threshold theorem, understanding how the endemic equilibrium depends smoothly on curvature, and extending the model to fractional, stochastic, or continuous spatial settings would significantly deepen the interplay between geometry and epidemiology.
Beyond theoretical interest, these findings provide structural insights relevant to empirical modelling of spatially heterogeneous diseases. Curvature captures patterns of communication, bottlenecks and anisotropies that traditional homogeneous weights cannot encode. As suggested by recent empirical analyses of dengue transmission in Recife, incorporating curvature into epidemic models may offer a principled mechanism for explaining spatial disparities, identifying persistent hotspots, and evaluating network-based interventions.
In this sense, curvature serves as a bridge between theory and application: a geometric parameter that not only shapes the mathematics of the model but is also measurable from real data. This dual character opens opportunities for future interdisciplinary work combining mathematical modelling, network geometry, spatial epidemiology and data science.
Overall, the framework developed in this article supports the view that epidemic dynamics on networks with intrinsic geometry can be governed by structural principles that are spectral, geometric and dynamical in nature. The further development of these principles may lead to a unified theory capable of capturing both the complexity of real-world epidemics and the mathematical regularities that underlie their global behaviour.